How much transformed free energy would be required to provide cells with 7.8mM ATP at 37 Celcius if both ADP and inorganic phosphate are at 0.3mM
1
Expert's answer
2019-09-18T06:45:36-0400

Transformed free energy required for the production of ATP can be estimated using the following equation:

ΔG = ΔGo + RT ln([ATP]/[ADP][Pi]),

where ΔG - free energy of the reaction, ΔGo - standard free energy, R - gas constant, T - temperature, [ATP] - ATP concentration, [ADP] - ADP concentration, [Pi] - phospate concentration.

As ΔGo = 30 kJ/mol, R = 37°C = 310.15 K, R = 8.314 × 10-3 kJ/mol·K:

ΔG = 30 kJ/mol + 8.314 × 10-3 kJ/mol·K × 310.15 K × ln(7.8 mM / 0.3 mM × 0.3 mM) = 30 kJ/mol + 8.314 × 10-3 kJ/mol·K × 310.15 K × 4.462 = 30 kJ/mol + 11.5 kJ/mol = 41.5 kJ/mol

Answer: +41.5 kJ/mol